# Metallicities of M Dwarf Planet Host Stars from Kepler, K2, and TESS observed by APOGEE: Trends with Exoplanetary Radii and Orbital Periods

### v2

@@ 3.3 Trends of Exoplanet Radii with Orbital Periods for Large Planets
The presence of hot Jupiters with Porb 4 days is well known for exoplanets orbiting solar-type stars since early exoplanet detections based on RV measurements. In fact, the ’pile-up’ of large exoplanets around Porb=3–5 days is also observed for exoplanets around F, G, and K stars. To make a comparison with what we find for the M dwarfs using more recent statistics, we compiled mean parameters for each exoplanet/stellar host from the NASA Exoplanet Archive separating the stellar hosts into spectral types F (6000 K T 7600 K), G (5200 K T 6000 K), and K (4000 K T 5200 K). We then derived the median percentiles of the orbital period distribution for exoplanets with R4 R⊕ of 4.5 days for F-type hosts, 5.0 days for G-type hosts, and 4.8 days for K-type hosts. These median orbital periods are generally in line and all fall roughly within the orbital period limits for our M-dwarf sample in Figure 7. However, significant differences exist between the orbital-period distributions, where exoplanets orbiting M dwarfs are found to be much more confined in period, with less scatter, than the ones for exoplanets around F, G, and K stars, which extend to much longer orbital periods. Large exoplanets with P 5 days were barely detected around M dwarfs (2 planets out of 27 planets with Rp 4 R⊕).
@@
In contrast to the observed concentration of large planets in our sample, it is clear that the population of small exoplanets in Figure 7 spans a much wider range in orbital periods, showing no evidence of a ”pile up” at any orbital period. Another feature from Figure 7 is that all multis in our sample, except one, have R 3 R⊕. This result is in agreement with what is found for solar-type in stars in the literature. Dong et al. (2018) studied FGK hosts of exoplanets with orbital periods of up to 10 days using LAMOST data. They found that all hot Jupiters (R10 R⊕), and most Hot Neptunes (2 RR6 Roplus) from their sample are single-detected exoplanets. Figure 1 of their work shows that this signature is even cleaner if the threshold of R4 R⊕ is used, as we do here. This may suggest that multi-planetary systems might be less efficient in producing large exoplanets compared to single systems, but of course, it is also the case that there could be more exoplanets in singles that remain undetected. Unfortunately, given the unavailability of APOGEE spectra in this study, we are not able to derive metallicities for all exoplanet hosts in Figure 7, but it is important to keep in mind that for our subsample of host stars/exoplanets with derived host metallicities (red circles in Figure 6), we find that the host stars of multis are generally more metal-poor compared to those of singles. This suggests that the metallicity and availability of solids in the protoplanetary disk may play a dominant role in the formation of large planets. We also note that the population of the largest exoplanets of multis has longer orbital periods than 3.5 days, with only four of these having shorter periods than 4.3 days.
@@ 4 Conclusions
A sample comprised of 45 M dwarfs, which host 65 exoplanets, was analyzed using high-resolution near-IR spectra from the SDSS APOGEE survey (Majewski et al., 2017) via spectrum synthesis computed in LTE. We used the Turbospectrum code (Plez, 2012) and MARCS model atmospheres (Gustafsson et al., 2008) to derive stellar metallicities ([M/H]) and oxygen abundances ([O/H]). In addition, in this study, we examined the distribution of exoplanetary radii for a larger sample composed of 246 exoplanets orbiting 188 M-dwarf stars. This sample includes the exoplanets studied in Wanderley et al. (2025), which analyzed the radius gap of small planets around M dwarfs, and in this study, an additional 28 large exoplanets were added to the sample.
@@
G 9-40 | 27.82 0.02 | 3423 | 4.89 | -0.25 | 5.36 | 8.46 | 6.97 | 0.30 | G 9-40 b | 0.340.03 | 1.900.21 | 5.75 |
GJ 3470 | 29.38 0.02 | 3650 | 4.72 | 0.13 | 1.00 | 8.70 | 5.65 | 0.48 | GJ 3470 b | 0.60.02 | 4.060.26 | 3.34 |
K2-25 | 44.62 0.06 | 3260 | 4.86 | 0.14 | 9.27 | 8.77 | 7.20 | 0.27 | K2-25 b | 1.160.06 | 3.190.36 | 3.48 |
TOI-1685 | 37.6 0.02 | 3568 | 4.76 | 0.13 | 4.45 | 8.77 | 5.88 | 0.45 | TOI-1685 b | 0.130.01 | 1.720.13 | 0.67 |
The Teff, , metallicity, and stellar radii uncertainties are respectively 100 K, 0.15 dex, 0.14 dex, and 0.03 R⊙. The complete table is available in electronic format.

### v3

@@ 3.3 Trends of Exoplanet Radii with Orbital Periods for Large Planets
The presence of hot Jupiters with Porb 4 days is well known for exoplanets orbiting solar-type stars since early exoplanet detections based on RV measurements. In fact, the “pile-up” of large exoplanets around Porb=3–5 days is also observed for exoplanets around F, G, and K stars. To make a comparison with what we find for the M dwarfs using more recent statistics, we compiled mean parameters for each exoplanet/stellar host from the NASA Exoplanet Archive separating the stellar hosts into spectral types F (6000 K T 7600 K), G (5200 K T 6000 K), and K (4000 K T 5200 K). We then derived the median percentiles of the orbital period distribution for exoplanets with R4 R⊕ of 4.5 days for F-type hosts, 5.0 days for G-type hosts, and 4.8 days for K-type hosts. These median orbital periods are generally in line and all fall roughly within the orbital period limits for our M-dwarf sample in Figure 7. However, significant differences exist between the orbital-period distributions, where exoplanets orbiting M dwarfs are found to be much more confined in period, with less scatter, than the ones for exoplanets around F, G, and K stars, which extend to much longer orbital periods. Large exoplanets with P 5 days were barely detected around M dwarfs (2 planets out of 27 planets with Rp 4 R⊕).
@@
In contrast to the observed concentration of large planets in our sample, it is clear that the population of small exoplanets in Figure 7 spans a much wider range in orbital periods, showing no evidence of a “pile up” at any orbital period. Another feature from Figure 7 is that all multis in our sample, except one, have R 3 R⊕. This result is in agreement with what is found for solar-type in stars in the literature. Dong et al. (2018) studied FGK hosts of exoplanets with orbital periods of up to 10 days using LAMOST data. They found that all hot Jupiters (R10 R⊕), and most Hot Neptunes (2 RR6 Roplus) from their sample are single-detected exoplanets. Figure 1 of their work shows that this signature is even cleaner if the threshold of R4 R⊕ is used, as we do here. This may suggest that multi-planetary systems might be less efficient in producing large exoplanets compared to single systems, but of course, it is also the case that there could be more exoplanets in singles that remain undetected. Unfortunately, given the unavailability of APOGEE spectra in this study, we are not able to derive metallicities for all exoplanet hosts in Figure 7, but it is important to keep in mind that for our subsample of host stars/exoplanets with derived host metallicities (red circles in Figure 6), we find that the host stars of multis are generally more metal-poor compared to those of singles. This suggests that the metallicity and availability of solids in the protoplanetary disk may play a dominant role in the formation of large planets. We also note that the population of the largest exoplanets of multis has longer orbital periods than 3.5 days, with only four of these having shorter periods than 4.3 days.
@@ 4 Conclusions
A sample comprised of 48 M dwarfs, was analyzed using high-resolution near-IR spectra from the SDSS APOGEE survey (Majewski et al., 2017) via spectrum synthesis computed in LTE. We used the Turbospectrum code (Plez, 2012) and MARCS model atmospheres (Gustafsson et al., 2008) to derive stellar metallicities ([M/H]) and oxygen abundances ([O/H]). In addition, in this study, we examined the distribution of exoplanetary radii for a larger sample composed of 246 exoplanets orbiting 188 M-dwarf stars. This sample includes the exoplanets studied in Wanderley et al. (2025), which analyzed the radius gap of small planets around M dwarfs, and in this study, an additional 28 exoplanets were added to the sample.
@@
| G 9-40 | 27.82 0.02 | 3423 | 4.89 | -0.25 | 5.36 | 8.46 | 6.97 | 0.300.01 | G 9-40 b | 0.340.03 | 1.900.12 | 5.75 |
| GJ 3470 | 29.38 0.02 | 3650 | 4.72 | 0.13 | 1.00 | 8.70 | 5.65 | 0.480.01 | GJ 3470 b | 0.600.02 | 4.060.14 | 3.34 |
| K2-25 | 44.62 0.06 | 3260 | 4.86 | 0.14 | 9.27 | 8.77 | 7.20 | 0.270.02 | K2-25 b | 1.160.06 | 3.190.21 | 3.48 |
| TOI-1685 | 37.6 0.02 | 3568 | 4.76 | 0.13 | 4.45 | 8.77 | 5.88 | 0.450.01 | TOI-1685 b | 0.130.01 | 1.720.08 | 0.67 |
The Teff, , and metallicity uncertainties are respectively 100 K, 0.20 dex, and 0.14 dex. The complete table is available in electronic format.